show that for any three vectors a,b,c
a×(b×c)+b×(c×a)+c×(a×b)=0
1
Expert's answer
2020-04-20T18:27:12-0400

"{\rm Consider}\, \\
\vec{a}=(a_x,a_y,a_z); \vec{b}=(b_x,b_y,b_z); \vec{c}=(c_x,c_y,c_z).\\
(\vec{b}\times\vec{c})=
\begin{vmatrix}
 \vec{i} & \vec{j} & \vec{k} \\
 b_x & b_y & b_z \\
c_x & c_y & c_z
\end{vmatrix} 
= \vec{i} (b_y c_z - b_z c_y )- \\
\vec{j} ( b_x c_z - b_z c_x ) + \vec{k} (b_x c_y -b_y c_x);\\
\vec{a}\times(\vec{b}\times\vec{c})=\\
\begin{vmatrix}
 \vec{i} & \vec{j} & \vec{k} \\
 a_x & a_y & a_z \\
(b_y c_z - b_z c_y )& ( b_z c_x - b_x c_z ) & (b_x c_y -b_y c_x)
\end{vmatrix} =\\
\vec{i}(a_y (b_x c_y -b_y c_x)-a_z( b_z c_x - b_x c_z ) )-\\
\vec{j}( a_x(b_x c_y -b_y c_x)-a_z(b_y c_z - b_z c_y ))+\\
\vec{k}( a_x( b_z c_x - b_x c_z )-a_y(b_y c_z - b_z c_y ));\\
(\vec{c}\times\vec{a})=
\begin{vmatrix}
 \vec{i} & \vec{j} & \vec{k} \\
 c_x & c_y & c_z \\
a_x & a_y & a_z
\end{vmatrix} 
= \vec{i} (c_y a_z - c_z a_y )- \\
\vec{j} ( c_x a_z - c_z a_x ) + \vec{k} (c_x a_y -c_y a_x);\\
\vec{b}\times(\vec{c}\times\vec{a})=\\
\begin{vmatrix}
 \vec{i} & \vec{j} & \vec{k} \\
 b_x & b_y & b_z \\
(c_y a_z - c_z a_y )& ( c_z a_x - c_x a_z ) & (c_x a_y -c_y a_x)
\end{vmatrix} =\\
\vec{i}(b_y (c_x a_y -c_y a_x)-b_z( c_z a_x - c_x a_z ) )-\\
\vec{j}( b_x(c_x a_y -c_y a_x)-b_z(c_y a_z - c_z a_y ))+\\
\vec{k}( b_x( c_z a_x - c_x a_z )-b_y(c_y a_z - c_z a_y ));\\
(\vec{a}\times\vec{b})=
\begin{vmatrix}
 \vec{i} & \vec{j} & \vec{k} \\
 a_x & a_y & a_z \\
b_x & b_y & b_z
\end{vmatrix} 
= \vec{i} (a_y b_z - a_z b_y )- \\
\vec{j} ( a_x b_z - a_z b_x ) + \vec{k} (a_x b_y -a_y b_x);\\
\vec{c}\times(\vec{a}\times\vec{b})=\\
\begin{vmatrix}
 \vec{i} & \vec{j} & \vec{k} \\
 c_x & c_y & c_z \\
(a_y b_z - a_z b_y )& ( a_z b_x - a_x b_z ) & (a_x b_y -a_y b_x)
\end{vmatrix} =\\
\vec{i}(c_y (a_x b_y -a_y b_x)-c_z( a_z b_x - a_x b_z ) )-\\
\vec{j}( c_x(a_x b_y -a_y b_x)-c_z(a_y b_z - a_z b_y ))+\\
\vec{k}( c_x( a_z b_x - a_x b_z )-c_y(a_y b_z - a_z b_y ));\\




\vec{a}\times(\vec{b}\times\vec{c})+\vec{b}\times(\vec{a}\times\vec{a})+\vec{c}\times(\vec{a}\times\vec{b})=\\
(a_y (b_x c_y -b_y c_x)-a_z( b_z c_x - b_x c_z ) )+\\
(b_y (c_x a_y -c_y a_x)-b_z( c_z a_x - c_x a_z ) )+\\
(c_y (a_x b_y -a_y b_x)-c_z( a_z b_x - a_x b_z ) )-\\
(( a_x(b_x c_y -b_y c_x)-a_z(b_y c_z - b_z c_y ))+\\
( b_x(c_x a_y -c_y a_x)-b_z(c_y a_z - c_z a_y ))+\\
( c_x(a_x b_y -a_y b_x)-c_z(a_y b_z - a_z b_y ))
)+\\
( a_x( b_z c_x - b_x c_z )-a_y(b_y c_z - b_z c_y ))+\\
( b_x( c_z a_x - c_x a_z )-b_y(c_y a_z - c_z a_y ))+\\
( c_x( a_z b_x - a_x b_z )-c_y(a_y b_z - a_z b_y ))=0"